# Temporally Consistent Factuality Probing for Large Language Models

### v1

@@ 1 Introduction
The emergence of large language models (LLMs) significantly contributes to achieving broader objectives of Artificial General Intelligence (AGI) in acquiring self-learning capabilities to solve complex tasks Ge et al. (2023). As the landscape of LLMs is continuously evolving, the capability to comprehend, process, and reason about time, known as temporal reasoning, has emerged as one of the cornerstones for advanced cognitive processing in LLMs. Temporal reasoning goes beyond greater recognition of dates and duration; it embodies the convoluted understanding of temporal dependencies, chronological sequences, and the nuanced, consistent relationship of events across time.
@@
Contributions. In short, we make the following contributions through this study***We will opensource the source code and dataset.:
@@ 3 TeCFaP Task Structure
The second metric is temporal-consistency. Given a pair of prefix-style paraphrases for a subject_relation filled with an identical key_object, an identical value_object should be generated. The metric estimate is binary (one or zero) for a given pair of such paraphrases if the model’s responses are identical or different. Forward and backward directions paraphrases contribute to forward and backward temporal-consitency, respectively.
Temporally consistent factuality. The third composite metric, temporally-consistent-factuality, is a stricter version of temporal-factuality, requiring a model to be consistent and factual across the temporal direction. It reports the temporal-factuality only if the responses are identical from all prefix-style paraphrases for a given subject_relation and key_object in a particular temporal direction. Paraphrases in forward and backward directions contribute to forward and backward temporally-consitent-factuality, respectively.
Other metrics. The quality of patterns is measured through temporal-succ_patt, which reports the percentage of patterns that return a correct value_object at least once in the entire probe. Also, to estimate the degree of temporal knowledge in models, we define temporal-succ_objs. Here, we report the percentages of value_objects, which are correctly generated at least once by the model. Further, we define temporal-know_cons and temporal-unk_cons as metrics to measure temporal-consistency of the fraction of patterns which generated correct value_object at least once and the fraction of patterns which never responded with a correct value_object for a subject_relation, respectively. All four metrics are then classified into forward and backward temporal directions.
@@ 4 Consistent-Time-Sensitive Learning Framework (CoTSeLF)
Motivation. Instruction tuning (IT) Zhang et al. (2023) is widely used as a low-cost, efficient modeling technique in developing domain or task-specific models. In recent IT advancements, a parameter-efficient fine-tuning-based (PEFT) Mangrulkar et al. (2022) method such as low-rank adaptation (LoRA) Hu et al. (2021) has shown tremendous potential to apply effective instruction tuning on LLMs for low-cost infrastructure settings. Moreover, RL combined with supervised fine-tuning has proven effective in improving the fine-tuned model in several domain-specific tasks, including temporal reasoning Tan et al. (2023). Furthermore, advances in multi-task learning led to several breakthroughs in modeling multiple objectives simultaneously Zhang and Yang (2022).
@@
Consistent Time-Sensitive Reinforcement Learning (CTSRL). Multiple value_objects for a given key_object are possible by excluding the temporal direction. We propose CTSRL to model consistent sensitivity towards time and overcoming the limitation of binary temporal character defined by TSRL Tan et al. (2023). To further fine-tune $\theta$, we apply CTSRL to model consistent temporal sensitivity by jointly modeling k1 and k2. The reward function is the linear combination of temporal and consistent sensitivities components. First, we define $\texttt{CTSRL}_{Discrete}$ in line with the original reward model of TSRL, where a positive discrete reward of value equal to one is given for correct prediction and zero for a negative answer. We also define $\alpha$ as the weight parameter for the consistence sensitivity reward component.
@@
For a given input $x$, the overall reward function $R_{d}(x)$ in the $\texttt{CTSRL}_{Discrete}$ setting is presented in Equation 1, where $R_{d}^{t}(x)$ and $R_{d}^{c}(x)$ are reward contribution for tasks k1 and k2, respectively. In Equation 2, $\lambda$ is used as a task indicator where $P^{\lambda}_{d}(x)$ is a positive reward score for consistence sensitivity component if $\lambda$ equals to $c$. Similarly, $N^{\lambda}_{d}(x)$ is a negative reward score. ${O_{g}}(\theta(x))$ and $O_{l}(x)$ are the generated value_object and ground-truth label value_object, respectively for given input $x$. In the case of $\texttt{CTSRL}_{Discrete}$, a positive reward score equal to one is assigned in case of correctly generated output and zero otherwise for both tasks k1 and k2. We further define another variant, $\texttt{CTSRL}_{Smooth}$, to model continuous and relative properties of time. The temporal sensitivity reward is a continuous function where a positive reward has a maximum function value equal to one. However, the negative reward is proportional to the relative distance of the predicted answer from the correct answer in the temporal axis. The goal is to penalize incorrect answers that are distant from the correct answer more severely than the incorrect predictions that are nearby on the temporal axis. There is no change in the reward component for consistence sensitivity except for releasing the constraint of the parameter $alpha$, indicating that $\texttt{CTSRL}_{Smooth}$ is an unweighted linear combination of continuous temporal and discrete consistence sensitivity reward components.
@@ 5 Experimental Results
Temporally consistent factuality on TeCFaP. We start with a comparison of various LLMs in zero-shot setting. The task is to correctly complete a sentence with the expected value_object given an instruction followed by an input, i.e., "complete the given sentence with the correct phrase: Meteora was released by Linkin Park immediately after". We observe the destitute performance of all the experimented LLMs over both temporal-factuality and temporally-consitent-factuality in the range of [$0.95\%-3.63\%$] and [$0\%-1.48\%$], respectively (Table 2). We notice that GPT-J and Falcon tend to be highly consistent in the range of [$41.52\%-64.87\%$] while being miserably factually incorrect compared to the LLaMA model in the same range of parameters size. Additionally, we notice that various families of LLMs behave differently regarding their sensitivity towards temporal direction. However, temporal direction and temporal consistency do not significantly correlate. Further, a preliminary evaluation of TeCFaP in zero-shot setting for GPT-4 OpenAI et al. (2023) with a minimal sample size yields temporal-factuality of $31.0\%$. (Appendix A.2.6).
@@
In Table 3, the additional context provided with an input improves temporal-factuality by $5.22$ percentage points for an IT model. The MT-IT model improves temporal-factuality, temporal-consistency and temporally-consitent-factuality by $7.7\%$, $6.4\%$ and $90.2\%$, respectively, compared to the IT model. We further observe improvements of $12.7\%$, $10.9\%$ and $90.4\%$ in temporal-factuality, temporal-consistency and temporally-consitent-factuality, respectively, with CoTSeLF over the baseline, indicating that improving a model’s temporal consistency also positively impacts its temporal factuality. However, the inaccessibility of GPT-4 architecture limits us to assess the efficacy of CoTSeLF’s on this model. (the probabilistic space evolution under CoTSeLF is presented in Appendix A.2.4).
@@ 7 Related Work
Consistency in knowledge bases was extensively studied, with developments around data and methods to benchmark the degree of inconsistencies Hansen and Jaumard (2000); Andersen and Pretolani (2001); Thimm (2013) across several tasks; QA Kassner et al. (2020), reading comprehension Antol et al. (2015); Rajpurkar et al. (2016); Ribeiro et al. (2019), summarizing Xie et al. (2021); Roit et al. (2023) and NLI Li et al. (2019).
Notable works in temporal information extraction include TimeBank Pustejovsky et al. (2003) and TimeEval Verhagen et al. (2010). The annotations defined in these datasets are primarily for time-event and events relationships such as before/after. The remarkable advancement in studying temporal reasoning over KGs led to the development of datasets such as TEQUILA Jia et al. (2018a), TimeQuestions Jia et al. (2021), and CronQuesions Saxena et al. (2021), probing KG’s response in ranking entities for a given temporal query. Subsequently, datasets like TEMPLAMA Dhingra et al. (2022) and StreamingQA Liška et al. (2022) were curated for temporal reasoning in LMs. The TSQA Chen et al. (2021) dataset has addressed the drawback of the limited time span of queries, but it defines only time-event relationships. TEMP-REASON Tan et al. (2023) captures multiple facets of temporal reasoning – time-time, time-event, and event-event relationships over a longer time span. However, it does not help evaluate consistent temporal reasoning that our dataset, TEMP-COFAC, addresses. TEMP-COFAC also overcomes a significant limitation of TEMP-REASON’s monotonous anchor queries (anchored only to one entity type; person name as subjects or objects) by providing a diverse set of anchor queries via sixty-six unique sub-rel-obj triplets formed around eleven entity-types. Additionally, it has a broader representation of eleven entity types compared to six in TEMP-REASON (Appendix A.1).
@@ 9 Limitations
On the other hand, the challenges related to computational assets to stack LLMs continue. Due to asset limitations, we may not utilize more extensive or commercially accessible LLMs to comprehensively evaluate on TeCFaP and approve on the off chance that the preferences of CoTSeLF are also advantageous to those models. Moreover, the smooth variant of CoTSeLF looks promising from its conception point of view but it could not outperform its discrete variant’s peer. We leave an in-depth ablation on smooth variants for future research.
@@ 10 Ethics Statement
The subjects and entities selected to be part of TEMP-COFAC are prone to unintended human biases during construction. The authors do not propagate any views/opinions, products, or representations of these subjects or entities in any form. The fact that women do not find representation during TEMP-COFAC annotations should be seen as a symptom of the gender disparity in research and innovation worldwide, but this is not the authors’ view. We support gender and racial equality in research and innovation with the utmost sincerity. Additionally, no generative AI-based content creation tools or applications were used to create this artifact, except for specialized support for spell checking, grammar correction, and paraphrasing.
@@ A.2.5 Significance Test for Alpha
A.2.6 GPT-4 Evaluation Results
This section presents the evaluation results for a prominent commercial LLM, GPT-4. We conduct this experiment with a limited randomly selected 100 samples from the TEMP-COFAC due to financial limitations. We set up the probe in zero-shot, where the task is to complete a sentence with the correct phrase. To this purpose, an instruction (Instruction: "complete the given sentence with the correct phrase") is provided along with the input to the model. In a zero-shot setting, the temporal-factuality stands at $31.00\%$. Trained with a trillion of parameters, in absolute terms, the temporal-factuality in GPT-4 is not very impressive, although it surpasses the relatively mid-sized LLMs such as LLaMA with a substantial margin in the zero-shot setting. An example of each of the positive and negative responses is mentioned in Table 8. CoTSeLF’s significant improvements over LLaMA[13B] do not exceed the standard performance of GPT-4. This is expected given the large differences in their world knowledge through the respective parameter sizes and an access to the open internet in case of GPT-4. Furthermore, we could not assess CoTSeLF’s efficacy on GPT-4 model due to the inaccessibility of its architecture for open fine-tuning.

### v2

@@ 1 Introduction
Large Language Models (LLMs) are pivotal in propelling the advancement of Artificial General Intelligence (AGI) by acquiring self-learning capabilities for complex tasks Ge et al. (2023). A key development within LLMs is the ability for temporal reasoning - comprehending, processing, and reasoning about time-related concepts, temporal dependencies, chronological sequences, and the nuanced, consistent temporal relationship of events. This ability is vital for a myriad of domain-specific tasks, including but not limited to summarizing timelines, tracking disease progression (medical), scheduling events (planning), managing contracts (legal), historical analysis (archaeology), and identifying tasks dependencies (project management).
@@
Contributions. In short, we make the following contributions through this study***Source code and dataset are available at https://github.com/ab-iitd/tecfap:
@@ 3 TeCFaP Task Structure
The second metric is temporal-consistency. Given a pair of prefix-style paraphrases for a subject_relation filled with an identical key_object, an identical value_object should be generated. The metric estimate is binary (one or zero) for a given pair of such paraphrases if the model’s responses are identical or different. Forward and backward directions paraphrases contribute to forward and backward temporal-consistency, respectively.
Temporally consistent factuality. The third composite metric, temporally-consistent-factuality, is a stricter version of temporal-factuality, requiring a model to be consistent and factual across the temporal direction. It reports the temporal-factuality only if the responses are identical from all prefix-style paraphrases for a given subject_relation and key_object in a particular temporal direction. Paraphrases in forward and backward directions contribute to forward and backward temporally-consistent-factuality, respectively.
Other metrics. The quality of patterns is measured using temporal-succ_patt, indicating the percentage of patterns yielding a correct value_object at least once during the probe. Furthermore, temporal-succ_objs is introduced to measure the model’s temporal knowledge by reporting the percentage of value_objects accurately generated at least once. Further, we define temporal-know_cons and temporal-unk_cons as metrics to measure temporal-consistency of the fraction of patterns which generated correct value_object at least once and the fraction of patterns which never responded with a correct value_object for a subject_relation, respectively. All four metrics are then classified into forward and backward temporal directions.
@@ 4 Consistent-Time-Sensitive Learning Framework (CoTSeLF)
Motivation. Instruction tuning (IT) Zhang et al. (2023) is a cost-effective, efficient technique for developing specialized models, as evidenced by the parameter-efficient fine-tuning (PEFT) Mangrulkar et al. (2022) approach such as low-rank adaptation (LoRA) Hu et al. (2021), which offers significant benefits for instruction tuning in LLMs within low-cost infrastructures. Additionally, the combination of RL and supervised fine-tuning enhances model performance in several domain-specific tasks, notably temporal reasoning Tan et al. (2023). Furthermore, advances in multi-task learning led to several breakthroughs in modeling multiple objectives simultaneously Zhang and Yang (2022).
@@
Consistent Time-Sensitive Reinforcement Learning (CTSRL). Multiple value_objects for a given key_object are possible by excluding the temporal direction. We introduce CTSRL, aimed at modeling consistent sensitivity towards time and overcoming the limitations of the binary temporal characterization inherent in TSRL Tan et al. (2023). CTSRL is employed to further fine-tune $\theta$ through the joint modeling of both k1 and k2. The reward mechanism is devised as a linear amalgamation of temporal and consistence sensitivities aspects. Initially, $\texttt{CTSRL}_{Discrete}$ is conceptualized in alignment with TSRL’s binary reward framework, awarding a positive discrete reward of one for accurate predictions and zero for incorrect responses. Additionally, $\alpha$ is defined to act as the weighting parameter for the consistence sensitivity reward component.
@@
For a given input $x$, the overall reward function $R_{d}(x)$ in the $\texttt{CTSRL}_{Discrete}$ setting is presented in Equation 1, where $R_{d}^{t}(x)$ and $R_{d}^{c}(x)$ are reward contribution for tasks k1 and k2, respectively. In Equation 2, $\lambda$ is used as a task indicator where $P^{\lambda}_{d}(x)$ is a positive reward score for consistence sensitivity component if $\lambda$ equals to $c$. Similarly, $N^{\lambda}_{d}(x)$ is a negative reward score. ${O_{g}}(\theta(x))$ and $O_{l}(x)$ are the generated value_object and ground-truth label value_object, respectively for given input $x$. In the case of $\texttt{CTSRL}_{Discrete}$, a positive reward score equal to one is assigned in case of correctly generated output and zero otherwise for both tasks k1 and k2.
We further define another variant, $\texttt{CTSRL}_{Smooth}$, to model continuous and relative properties of time. The temporal sensitivity reward is a continuous function where a positive reward has a maximum function value equal to one. However, the negative reward is proportional to the relative distance of the predicted answer from the correct answer in the temporal axis. The goal is to penalize incorrect answers that are distant from the correct answer more severely than the incorrect predictions that are nearby on the temporal axis. There is no change in the reward component for consistence sensitivity except for releasing the constraint of the parameter $alpha$, indicating that $\texttt{CTSRL}_{Smooth}$ is an unweighted linear combination of continuous temporal and discrete consistence sensitivity reward components.
@@ 5 Experimental Results
Temporally consistent factuality on TeCFaP. We start with a comparison of various LLMs in zero-shot setting. The task is to correctly complete a sentence with the expected value_object given an instruction followed by an input, i.e., "complete the given sentence with the correct phrase: Meteora was released by Linkin Park immediately after". We observe the destitute performance of all the experimented LLMs over both temporal-factuality and temporally-consitent-factuality in the range of [$0.95\%-3.63\%$] and [$0\%-1.48\%$], respectively (Table 2). We notice that GPT-J and Falcon tend to be highly consistent in the range of [$41.52\%-64.87\%$] while being miserably factually incorrect compared to the LLaMA model in the same range of parameters size. Additionally, various families of LLMs behave differently regarding their sensitivity towards temporal direction, but it doesn’t significantly correlate with temporal consistency. Further, a preliminary evaluation of TeCFaP in zero-shot setting for commercial LLMs such as GPT-4 OpenAI et al. (2023) and Claude-3 is Anthropic (2024) presented in Appendix A.4.
@@
In Table 3, the additional context provided with an input improves temporal-factuality by $5.22$ percentage points for an IT model. The MT-IT model improves temporal-factuality, temporal-consistency and temporally-consitent-factuality by $7.7\%$, $6.4\%$ and $90.2\%$, respectively, compared to the IT model. We further observe improvements of $12.7\%$, $10.9\%$ and $90.4\%$ in temporal-factuality, temporal-consistency and temporally-consitent-factuality, respectively, with CoTSeLF over the baseline, indicating that improving a model’s temporal consistency also positively impacts its temporal factuality. However, the inaccessibility of GPT-4 architecture limits us to assess the efficacy of CoTSeLF’s on this model. (the probabilistic space evolution under CoTSeLF including the ablations for scalability are presented in Appendix A.2.4, & A.3).
Significance of CTSRL over TSRL. We perform this ablation in two different scenarios by immobilizing the SFT model. In the first scenario, an IT model serves as the foundation [comparing (IT + TSRL) with (IT + $\texttt{CTSRL}_{Discrete}$)], while in the second scenario, the basis is the MT-IT model [comparing (MT-IT + TSRL) with (MT-IT + $\texttt{CTSRL}_{Discrete}$)]. These experiments are meticulously executed under identical conditions, as meticulously outlined in Table 3.
Model Temp-fact Temp-cons Temp-cons-fact
Base SFT Model: IT
TSRL 16.60 33.23 2.28
CTSRL 17.6 34.27 2.74
Base SFT Model: MT-IT
TSRL 17.55 34.5 3.68
CTSRL 18.72 36.88 4.34
Table 4: Comparison of $\texttt{CTSRL}_{Discrete}$ and TSRL by immobilizing the SFT model across two settings: IT and MT-IT model.
The findings are delineated in Table 4. $\texttt{CTSRL}_{Discrete}$ exhibits superior performance over TSRL on all metrics: temporal-factuality, temporal-consistency, and temporally-consistent-factuality in both experimental settings. Results from this ablation study further underscore the distinct advantage of a preference of CTSRL over TSRL.
@@ 7 Related Work
A noteworthy advancement in temporal factuality involves analyzing and updating LLMs to address the obsolescence of their factual knowledge over time Hu et al. (2024); Vu et al. (2023). These approaches concentrate on amending the factuality of evolving temporal relationships without rectifying the current inaccuracies and inconsistencies. On the other hand, consistency in KBs was extensively studied, with developments around data and methods to benchmark the degree of inconsistencies Hansen and Jaumard (2000); Andersen and Pretolani (2001); Thimm (2013) across several tasks; QA Kassner et al. (2020), reading comprehension Antol et al. (2015); Rajpurkar et al. (2016); Ribeiro et al. (2019), summarizing Xie et al. (2021); Roit et al. (2023) and NLI Li et al. (2019).
Notable works in temporal information extraction include TimeBank Pustejovsky et al. (2003) and TimeEval Verhagen et al. (2010). The annotations defined in these datasets are primarily for time-event and events relationships such as before/after. The remarkable advancement in studying temporal reasoning over KGs led to the development of datasets such as TEQUILA Jia et al. (2018a), TimeQuestions Jia et al. (2021), and CronQuesions Saxena et al. (2021), probing KG’s response in ranking entities for a given temporal query. Subsequently, datasets like TEMPLAMA Dhingra et al. (2022) and StreamingQA Liška et al. (2022) were curated for temporal reasoning in LMs. The TSQA Chen et al. (2021) dataset has addressed the drawback of the limited time span of queries, but it defines only time-event relationships. TEMP-REASON Tan et al. (2023) captures multiple facets of temporal reasoning – time-time, time-event, and event-event relationships over a longer time span. Our dataset, TEMP-COFAC, addresses the lack of consistent temporal reasoning evaluation and improves upon TEMP-REASON by offering a diverse set of anchor queries through sixty-six unique sub-rel-obj triplets across eleven entity types, overcoming TEMP-REASON’s limitation of single-entity-type anchors and coverage of only six entity types. (Appendix A.1).
@@ 9 Limitations
Moreover, the underlying rationale for the smooth variant of CTSRL posits that a fundamental characteristic of time is its relativity. The objective is to impose greater penalties on incorrect responses that significantly deviate from the correct answer compared to those incorrect predictions that are proximal on the temporal scale. This premise suggests that such a formulation will compel the model to assimilate the relative aspect of time, thereby enhancing its efficiency in addressing queries necessitating relative temporal responses (before/after). While the smooth variant of CoTSeLF appears promising at the conceptual level, it has not surpassed the performance of its discrete counterparts. A comprehensive ablation study on smooth variants is earmarked for future investigation.
While temporal relation settings such as t+/-1 scenarios necessitate precision, as only a single correct response is viable for any given query contrary to the open temporal relation settings (one-to-many), it is our fervent hope that the insights derived from this study will inspire further research to test TeCFaP toward addressing multi-hop temporal queries (t+/-n) in forthcoming endeavors. On the other hand, the challenges related to computational assets to stack LLMs continue. Due to asset limitations, we may not utilize more extensive or commercially accessible LLMs to comprehensively evaluate on TeCFaP and approve on the off chance that the preferences of CoTSeLF are also advantageous to those models.
@@ 10 Ethics Statement
The subjects and entities selected to be part of TEMP-COFAC are prone to unintended human biases during construction. The authors do not propagate any views/opinions, products, or representations of these subjects or entities in any form. The fact that women do not find representation during TEMP-COFAC annotations should be seen as a symptom of the gender disparity in research and innovation worldwide, but this is not the authors’ view. We support gender and racial equality in research and innovation with the utmost sincerity. Furthermore, no generative AI-based content creation tools or applications were used to create this artifact, except for specialized support for spell checking, grammar correction, and paraphrasing.
Acknowledgement
Tanmoy Chakraborty acknowledges the support of the IBM-IITD AI Horizons network.
@@ A.2.5 Significance Test for Alpha
Model [Parameter Size] [Quantization] Temp-fact Temp-cons Temp-cons-fact
LLaMA[7B][8-bit] 8.10 24.07 0.00
LLaMA[13B][8-bit] 18.72 36.88 4.34
LLaMA[30B][4-bit] 23.05 46.55 6.04
Table 9: CoTSeLF’s ($\texttt{CTSRL}_{Discrete}$) results across various parameter sizes of LLaMA.
A.2.6 Significance Test for Sample Size
in Section 5, we analyzed sampling efficiency across temporal variations in entity size and their corresponding performance, as depicted in Figure 5. Furthermore, an additional experiment was undertaken to examine the relationship between the distribution of entity types and their respective performance within the test data set. Our findings indicate a notable negative correlation, quantified as -0.61, between the distribution of entity types and temporal-factuality. Likewise, a negative correlation, valued at -0.42, was observed in relation to the distribution of entity types and temporally-consistent-factuality. It was discerned that performance pertaining to domain-specific entity types, such as Software, Satellites, and Games, surpassed that of more general entity types, like Location and Person, with the data distribution being inversely skewed towards the latter.
A.3 Discussion on CoTSeLF’s Scalability
In the majority of instances, the solutions that rely on parameters fine-tuning, regardless of the specific large language model (LLM) architecture involved, demonstrate scalability and adaptability to diverse models without being constrained by the number of parameters. This ablation study is performed to investigate the scalability of the proposed solution, CoTSeLF, specifically its capacity to scale according to the model parameter size. For this purpose, the CoTSeLF is implemented in two different LLMs configurations, one with a reduced parameter size, LLaMA[7B]  and another with an increased parameter size, LLaMA[30B]. However, it was necessary to adjust the quantization from 8-bit to 4-bit for LLaMA[30B] and MT-IT’s LoRA adapter both due to computational resource constraints. The experiment is conducted in same conditions as outlined in Table 3
The findings from Table 9 reaffirm CoTSeLF’s efficacy for models across a spectrum of parameter sizes, inclusive of those with large dimensions. A pronounced correlation between CoTSeLF’s performance and the sizes of the parameters was noted. Moreover, it was observed that enhancements attributed to CoTSeLF are significant across all evaluated metrics with LLaMA[30B] despite an increment in quantization level.
A.4 Discussion on Commercial LLMs
This section presents the evaluation results for two prominent commercial LLMs, GPT-4 and Claude-3. We conducted this experiment with a limited randomly selected 100 samples from the TEMP-COFAC test set due to financial limitations. We set up the probe in zero-shot, where the task is to complete a sentence with the correct phrase. To this purpose, an instruction (Instruction: "complete the given sentence with the correct phrase") is provided along with the input to the model. Subsequently, we provide a concise examination of the potential for employing CoTSeLF within the realm of commercial LLMs applications.
A.4.1 GPT-4 Evaluation Results
We use ’GPT-4-0125-preview’ version of GPT-4 for this experiment. In a zero-shot setting, the temporal-factuality stands at $36.00\%$ (Table 10). Trained with a trillion of parameters, in absolute terms, the temporal-factuality in GPT-4 is not very impressive, and requires interventions such as CoTSeLF to improve it. An example of each of the positive and negative responses is mentioned in Table 11.
A.4.2 Claude-3 Evaluation Results
In conducting a parallel investigation with another leading commercial LLM, Claude-3, under identical conditions, we operationalized the ’CLAUDE-3-OPUS-20240229’ version for this assessment. The temporal-factuality of Claude-3 is noted at $20.00\%$ (Table 10), marking a notable decline in performance relative to GPT-4.